OpenStudy (anonymous):

We are given the inverse variation equation xy = ¼. Complete the following statement: x ________ as y decreases. increases decreases remains constant

OpenStudy (anonymous):

increases, example x = 1/4 y = 1 x = 1, y = 1/4 x = 2 y = 1/8